Pictorally with explanations, illustrate the necessity of using common denominators when adding fractions with different denominators. (I am in school for being a teacher and this concept seems hard to teach)
Answer by stanbon(57219)

Pictorally with explanations, illustrate the necessity of using common denominators when adding fractions with different denominators
-----------------------------------------------------------------
Let say you want to add 1/2 and 1/3.
------------
Draw two pies.
Split the 1st pie into two parts.
Split the 2nd pie into three parts.
---------------------------
How would you add 1 part of the two to 1 part of the three?
----------------------------
Hmmmmmm???
----------------------------
Color 1 of the 2 parts; Color 1 part of the three
---------------------------------------------------
Split each part of the 1st pie into three parts; you now have 6 parts.
Split each part of the 2nd pie into two parts; you now have 6 parts
----------------------------------------------------
Now you can add the 3 colored parts of the 1st pie to 2 colored parts
of the 2nd pie to be 5 parts out of 6.
------------------------------------------
(1/2) + (1/3) = (3/6) + (2/6) = 5/6
=============================================
Cheers,
Stan H.
-----------------------------